OpenStudy (anonymous):

The tangent line approximation for f(x)=sqrt (x^2+16) near x=-3 is...?

OpenStudy (anonymous):

@amistre64 @ganeshie8

OpenStudy (anonymous):

@zepdrix @hartnn

hartnn (hartnn):

so what are the steps to do a tangent line approx ? you know ?

OpenStudy (anonymous):

it is the same as local linearization i think?

hartnn (hartnn):

yes \(f(x) \approx f(a) + f'(a) (x-a)\)

hartnn (hartnn):

a =-3 here

OpenStudy (anonymous):

ok. so a=-3, so f(x)=f(-3)+f'(-3)(x+3)

OpenStudy (anonymous):

and \[f'(x)=\frac{ 2x }{ 2\sqrt{x+16} }\]

hartnn (hartnn):

yes

hartnn (hartnn):

if you mean x^2 +16

OpenStudy (anonymous):

\[f(x)= \sqrt{7}-\frac{ 3\sqrt{7} }{ 8}(x+3)\]

hartnn (hartnn):

sqrt 7 ? 8? howw ?

OpenStudy (anonymous):

i think my math's wrong, gimme a sec

hartnn (hartnn):

ohh (-3)^2 is not -9 its +9

OpenStudy (anonymous):

oh uhggggggg I did -3^2 is -9

hartnn (hartnn):

try again, you will get it this time :)

OpenStudy (anonymous):

\[f(x)=5-\frac{ 3 }{ \sqrt{13} }(x+3)\]

OpenStudy (anonymous):

I think that's right

hartnn (hartnn):

13 :O ??

hartnn (hartnn):

from where did 13 come ?

OpenStudy (anonymous):

must I rationalize denominator?

OpenStudy (anonymous):

oh my derivative is wrong

OpenStudy (anonymous):

sorry I'm terrible at this :/

hartnn (hartnn):

derivative is correct, plugging in might not be correct

hartnn (hartnn):

oh and as i told its x^2, not x

hartnn (hartnn):

so x^2 +16 = 9 +16 = 25 not -3+16

OpenStudy (anonymous):

\[f(x)= 5- \frac{ 3 }{ 5} (x+3)\]

hartnn (hartnn):

yes, there you go!

OpenStudy (anonymous):

thank you so much! I was just making silly math errors.

hartnn (hartnn):

just a moment

hartnn (hartnn):

why is there a negative ?

OpenStudy (anonymous):

a=-3, plugging into \[\frac{ x }{ \sqrt{x ^{2}-16} }\] gives you -3/5

hartnn (hartnn):

sorry :P thats correct

OpenStudy (anonymous):

:)